# Interpretability Illusions with Sparse Autoencoders: Evaluating Robustness of Concept Representations

### v1

@@ Abstract
Sparse autoencoders (SAEs) are commonly used to interpret the internal activations of large language models (LLMs) by mapping them to human-interpretable concept representations. While existing evaluations of SAEs focus on metrics such as the reconstruction-sparsity tradeoff, human (auto-)interpretability, and feature disentanglement, they overlook a critical aspect: the robustness of concept representations to input perturbations. We argue that robustness must be a fundamental consideration for concept representations, reflecting the fidelity of concept labeling. To this end, we formulate robustness quantification as input-space optimization problems and develop a comprehensive evaluation framework featuring realistic scenarios in which adversarial perturbations are crafted to manipulate SAE representations. Empirically, we find that tiny adversarial input perturbations can effectively manipulate concept-based interpretations in most scenarios without notably affecting the outputs of the base LLMs themselves. Overall, our results suggest that SAE concept representations are fragile and may be ill-suited for applications in model monitoring and oversight. The code for our experiments is available at https://github.com/AI4LIFE-GROUP/sae_robustness.
1.  Introduction
@@
In this work, we propose incorporating robustness as a core component of the SAE evaluation agenda, and concept-based explanations in general. We begin by defining SAE robustness under input-level perturbations, followed by the specification of adversarial settings for a comprehensive robustness evaluation. Each setting is characterized by two key objectives and a perturbation mode, which determines how the input-level attack is applied. First, the semantic goal: the perturbation could be either untargeted, aiming to alter the SAE activations from the original, or targeted, aiming to match the perturbed activations to that of another semantically unrelated input. Second, the activation goal: either at the population level, where the objective is to manipulate all SAE neurons simultaneously to alter the overall activation pattern; or at the individual level, where the goal is to activate a specific feature independently. Additionally, we consider two basic attack modes in this work: suffix-based attacks, where new tokens are appended to the input, and replacement-based attacks, where original tokens are allowed to be substituted with adversarial ones. Putting these together, we define eight different scenarios to evaluate SAE concept robustness against adversarial perturbations.
To search for adversarial inputs that manipulate SAE interpretations, we adapt Greedy Coordinate Gradient (GCG) (Zou et al., 2023), a widely used input-space search algorithm originally developed to generate adversarial prompts that induce harmful outputs in LLMs, to our SAE settings. Our findings reveal that current SAEs are not sufficiently robust to serve as reliable tools for interpreting and monitoring LLM behaviors, nor as trustworthy components in downstream applications that depend on stable, concept-level representations. We further believe that this vulnerability to adversarial input manipulation could generalize to other SAE-inspired approaches such as transcoders (Dunefsky et al., 2024) and crosscoders (Lindsey et al., 2024), highlighting the need for future research on developing more robust and reliable concept-extraction methods in the field of mechanistic interpretability.
@@
  • We propose a comprehensive evaluation framework that defines SAE robustness along semantic goals, activation goals, and perturbation modes, resulting in eight distinct evaluation scenarios.
  • We conduct extensive experiments by designing adversarial input-level attacks, showing that SAE interpretations are consistently vulnerable to input perturbations across different datasets and model configurations.
2.  Related Work
@@
3.  Evaluating the Robustness of SAE Interpretations
@@ 3.2 Proposed Theory of SAE Robustness
$\exists c_{1},c_{1}^{\prime}\in\mathcal{C}:d_{c}(c_{1},c_{1}^{\prime})<\epsilon_{c},d_{z}(z_{1},z_{1}^{\prime})>\delta_{z}$ (3)
$\exists c_{1},c_{2}\in\mathcal{C}:d_{c}(c_{1},c_{2})>\delta_{c},d_{z}(z_{1},z_{2})<\epsilon_{z}$ (4)
@@
  • They support more fine-grained control, allowing perturbations in arbitrary directions and with variable step sizes in the input space (as explained in Section 3.3).
@@
$\max_{x_{1}^{\prime}}d_{z}(z_{1},z_{1}^{\prime})\quad\text{subject to}\quad d_{x}(x_{1},x_{1}^{\prime})\leq\epsilon_{x}$ (5)
$\min_{x_{2}}d_{z}(z_{1},z_{2})\quad\text{subject to}\quad d_{x}(x_{1},x_{2})\geq\delta_{x}$ (6)
@@ 3.3.1 Semantic Goal
$\min_{x_{1}^{\prime}}d_{z}(x_{1}^{\prime},x_{2})\quad\text{subject to}\quad d_{x}(x_{1},x_{1}^{\prime})\leq\epsilon_{x}$ (7)
@@ 3.3.2 Activation Goal
The goal is to manipulate groups of SAE features simultaneously to control the overall sparse representation vector, so the distance metric can be defined by the overlap ratio between two sets of $k$ most activated SAE latents (we call this metric neuron overlap ratio throughout this work):
@@
The perturbation may also be directed toward a specific SAE feature by modifying its rank among all latent dimensions. In the case of activation, the objective is to increase the feature’s rank until it has non-zero activation after $\phi$. Conversely, in the case of deactivation, the objective is to reduce the feature’s rank such that the latent becomes zero. We thus define $d_{z}$ in a binary manner:
@@
These two levels capture complementary structural vulnerabilities of SAE representations. The population-level objective assesses global interpretation shifts, while the individual-level objective probes the microscopic stability of specific SAE features.
@@ 3.3.3 Perturbation Mode
Analogous to adversarial attacks on LLMs, the perturbation mode defines how adversarial edits are applied to the input. In this work, we consider two fundamental types:
Suffix mode
Adversarial tokens are appended to the end of the original input. This preserves the original semantic content while allowing the attacker to introduce new features that can dominate the existing ones by appending a limited-length suffix.
Replacement mode
Tokens in the original input are substituted with adversarial ones. Compared to the suffix mode, this approach directly modifies existing semantic features, and we therefore restrict the perturbation to a single token replacement to ensure that the change at the input level remains minimal. The overall robustness under replacement mode can be measured by the average across all token positions.
These two modes represent different types of adversarial control, and both are highly plausible in practical deployment scenarios. By systematically combining semantic goal, activation goal, and perturbation mode, we have defined a total of eight different settings to evaluate SAE robustness.
@@ 3.4 Generalized Input-level Attack for SAE
The complete pseudocode for our generalized input-level attack is provided in Appendix A.
4.  Experimental Evaluation
In this section, we first describe our experimental setup, including model and dataset choices, followed by results across all eight adversarial scenarios. We then present additional analyses to validate the reliability of our evaluation framework, examine robustness trends across model depths, assess cross-model attack transferability, and conclude with a case study on manipulating highly interpretable SAE latents.
4.1 Experimental Setup
Models
We evaluate SAE robustness on two combinations of open-source LLMs and pretrained SAEs on the residual streams of the models. In this section, we present the results for (1) layer 20 of Llama-3-8B with TopK as the SAE activation function (Gao et al., 2024) and (2) layer 30 of Gemma-2-9B with JumpReLU SAE (Rajamanoharan et al., 2024b). Both SAEs have a width of 131k. Our experiments are not conducted on smaller models or SAEs with fewer latents, as our targeted attacks require a relatively large initial distance $d_{z}(x_{1},x_{2})$ to meaningfully evaluate their effectiveness. In other words, we evaluate the robustness of an SAE only when it can clearly distinguish semantically unrelated inputs. We focus on mid-to-late layers, as they strike a balance between low-level surface patterns in early layers and highly entangled representations in final layers, making them better suited for capturing human-interpretable concepts. Additional results for different model depths are included in Appendix F.
Datasets
The ideal evaluation for targeted tasks should be conducted on two sets of input prompts with minimal semantic overlap. Our first ChatGPT-generated Art & Science dataset consists of 100 pairs of sentences each with approximately 20 to 25 tokens, broadly divided into two categories: $x_{1}$ is related to art and humanity, and $x_{2}$ is about science and technology. We adapt another existing dataset AG News to sample 300 pairs of news titles with different category labels as an additional evaluation dataset. The mean initial overlap ratios are $31.3\%$ and $33.4\%$ for the two datasets, averaged across two models.
Evaluation Configurations
Since the residual streams of LLMs encode both semantic features and next-token prediction information, we append a short instruction prompt to the original sequence (including the suffix), "The previous sentence is about", to the input to better extract LLM’s semantic content from the last hidden state. At individual level, we focus on $10$ individual neurons, which are selected for a activation/deactivation task based on the semantic goal: for an untargeted task, we select the SAE latents with the lowest/highest activation values for $x_{1}$; for a targeted task, we select the latents most/least activated by $x_{2}$ but are currently deactivated/activated by $x_{1}$. The specific choices of hyperparameters are included in Appendix B.
Model Semantic Goal Population Level Individual Level
Neuron Overlap (%) Attack Success Rate (%)
Suffix Replacement Suffix Replacement
Llama 3 8B Untargeted $89.1\pm 2.3$ $84.6\pm 0.9$ $\pm 0.0,22.9\pm 1.5 0.0±0.0$$\pm 0.0,34.3\pm 0.9$
Targeted $74.5\pm 3.5 +$$29.2\pm 1.9$ $55.9\pm 2.2,91.4\pm 0.8$ $59.2\pm 1.1,86.4\pm 0.8$
Gemma 2 9B Untargeted $85.5\pm 1.9$ $79.8\pm 1.1$ $\pm 0.0,30.5\pm 1.7 0.0±0.0$$\pm 0.0,32.6\pm 1.2$
Targeted $68.8\pm 3.8 +$$31.9\pm 2.2$ $63.0\pm 1.9,82.3\pm 1.4$ $61.3\pm 1.2,84.8\pm 0.6$
Table 2: Attack results for all eight settings for Llama-3-8B and Gemma-2-9B on our generated prompt dataset. For individual level, we include results for both activation (left) and deactivation (right) tasks. The mean values and standard deviations are computed based on three independent runs.
Figure 1: Examples of GCG and evaluation losses across iterations for targeted (left) and untargeted (right) population-level suffix attacks.
Example: Targeted Population-level Suffix Attack $x_{1}$: His research surveys Islamic architecture from Morocco to India, revealing regional adaptations and shared aesthetic principles $x_{2}$: Autonomous vehicles rely on sensor fusion, machine learning, and path planning to navigate complex traffic environments $x^{\prime}_{1}$: His research surveys Islamic architecture from Morocco to India, revealing regional adaptations and shared aesthetic principles UAV AMC algorithms
Example: Untargeted Population-level Suffix Attack $x_{1}$: Philosophical skepticism challenges empirical certainty, questioning whether knowledge is possible in the face of doubt $x^{\prime}_{1}$: Philosophical skepticism challenges empirical certainty, questioning whether knowledge is possible in the face of doubt setImageBitmap
Figure 2: Instances of adversarial attacks correspond to the loss curves in Figure 1. Adversarial suffixes are highlighted in red.
4.2 Results Interpretation
We include the complete results for our generated Art & Science dataset in Table 2, and provide the loss curves and adversarial inputs for two representative attack examples in Figure 1 and 2. For tasks on population level, we report the change in overlap ratio (as a percentage) between two sets of $k$ most activated latents, while for those on individual level we report the attack success rate (ASR). The experiment details and results for the AG News dataset are included in Appendix C. In general, our attacks are effective in most cases, and we discuss several important insights below.
Failure Cases
The adversarial attack is only ineffective in the untargeted tasks at the individual level: the SAE latents with lowest activation values cannot be activated by input perturbations. This is most likely due to dead latents (i.e. latents cannot be activated under any circumstances) prevalent in pretrained SAEs, which is aligned with previous findings by Gao et al. (2024).
Untargeted vs. Targeted
Suffix vs. Replacement
At the population level, suffix mode is more effective at controlling the overall SAE activation pattern than replacement mode. This trend is probably due to (1) more recent tokens are likely to have greater impact on both the residual stream and the SAE activations, and (2) appending a short suffix provides more fine-grained and predictable optimization steps on the latent representation space. When perturbing individual SAE features, the effectiveness of two perturbation modes varies across semantic and activation goals, with no significant implications.
4.3 Additional Analyses
Decoupling SAE and LLM Robustness
A potential critique of our evaluation strategy is that it does not fully disentangle the robustness of the SAE from that of the underlying LLM. To address this, we select the number of manipulatable tokens (e.g. suffix length) by examining the LLM’s generations before and after adversarial attacks. Since our user prompts end with an instruction to summarize the preceding sentence (see Section 4.1), the generated outputs reflect the LLM’s self-summarization of the semantic content. Then we can leverage an LLM judge (GPT-4.1-mini) to determine whether the semantic content changes notably after the attacks. We evaluate 100 successful attacks for each unique level of input manipulation, and when the total number of attacks exceed 100, we sample randomly while preserving the balance between different settings and token indices (for replacement mode). For each attack, we take the majority vote among 5 generations as the final decision. Based on the results for the suffix lengths and single-token replacement used in our experiments reported in Table 3, we show that our adversarial attacks are mostly targeting the SAEs, without notably affecting the base LLM. The prompt template and self-summarization examples are included in Appendix E.
Model Dataset Suffix (1 token) Suffix (3 tokens) Replacement (1 token)
Llama 3 8B Art & Science $0.98$ $0.93$ $1.0$
AG News $0.95$ - $0.96$
Gemma 2 9B Art & Science $1.0$ $0.94$ $1.0$
AG News $0.94$ - $0.95$
Table 3: Fraction of adversarial attacks that result in no significant differences in model’s self-summarization, using LLM-as-a-judge (GPT-4.1-mini). During generation, the temperature is set to 0.7 and top_p is set to 0.9.
Robustness Change Across Model Depth
The main results presented in this section focus on a single mid-to-late layer for each model. To assess the generalizability of our findings across different model depths, we apply our population-level attacks under suffix mode to additional layers using open-source pretrained SAE weights. Empirically, we find that our attacks remain effective across layers, despite slight decreases in performance upper bound. The experiment results and interpretations are included in Appendix F.
Transferability of Attacks
In reality, the adversary might want to craft a single perturbed sequence that could be used to attack different models, so it’s important to investigate the transferability of our attacks across different LLMs. Since SAE latents of different models encode different semantic concepts, we only investigate transferability at the population level. As presented in Appendix G, we observe that the difference between average neuron overlap changes achieved by the original attacks and transferred attacks are at most $25.3\%$ for the targeted settings and $43.8\%$ for the untargeted settings, suggesting that although there’s a notable performance gap due to the model transfer, the attacks are still effective.
Deactivating Highly Interpretable SAE Latents from Neuronpedia
In the above experiments, the set of manipulated SAE latents is empirically determined by our datasets (i.e., the most or least activated ones by our inputs). In fact, such evaluation procedure could be reversed: we may instead begin by selecting specific target SAE latents and then assess their robustness by identifying input sequences that strongly activate them. To explore this, we select several SAE latents associated with consistent and human-interpretable semantic concepts from Neuronpedia, and apply the untargeted individual-level attack to top-activating sequences drawn from an external text corpus. Through experiments, we find that our attack could successfully deactivate these highly meaningful SAE latents, with illustrative examples provided in Appendix H.
5.  Discussion and Limitations
As discussed in section 4.1, we target medium-sized LLMs because smaller models, along with their SAEs, lack the capacity to distinguish semantically unrelated inputs. However, this choice results in slower attacks: for instance, a 50-iteration targeted population-level suffix attack on a longer prompt from the Art & Science dataset takes approximately 10 minutes. In practical applications, users may only need to evaluate the most relevant scenario among our eight settings. Additionally, future work could explore optimizing the attack pipeline to better balance effectiveness and efficiency.
More broadly, we consider our evaluation framework and adversarial attacks as a general methodology for assessing concept-extraction tools for LLMs. While we focus on standard SAEs in this work, the same vulnerabilities likely extend to other variants such as transcoders (Dunefsky et al., 2024) and crosscoders (Lindsey et al., 2024), which similarly lack structural constraints or robustness-aware objectives during training. We leave the exploration of such extensions to future work.
6.  Conclusion
In this work, we investigate the robustness of SAEs under input-level adversarial perturbations and introduce a comprehensive evaluation framework spanning semantic and activation-level objectives. Our experiments show that SAE interpretations are highly vulnerable to minimal input changes, even when the underlying LLM remains semantically stable, raising concerns about their reliability in realistic settings. To advance trustworthy interpretability, we hope our work motivates the development of more robust tools for understanding LLMs, as stability under real-world conditions is essential for aligning model behavior with human expectations.
Appendix A Pseudocode for Adversarial Input-Level Attack

### v2

# Evaluating Adversarial Robustness of Concept Representations in Sparse Autoencoders

@@ Abstract
Sparse autoencoders (SAEs) are commonly used to interpret the internal activations of large language models (LLMs) by mapping them to human-interpretable concept representations. While existing evaluations of SAEs focus on metrics such as the reconstruction-sparsity tradeoff, human (auto-)interpretability, and feature disentanglement, they overlook a critical aspect: the robustness of concept representations to input perturbations. We argue that robustness must be a fundamental consideration for concept representations, reflecting the fidelity of concept labeling. To this end, we formulate robustness quantification as input-space optimization problems and develop a comprehensive evaluation framework featuring realistic scenarios in which adversarial perturbations are crafted to manipulate SAE representations. Empirically, we find that tiny adversarial input perturbations can effectively manipulate concept-based interpretations in most scenarios without notably affecting the base LLM’s activations. Overall, our results suggest that SAE concept representations are fragile and without further denoising or postprocessing they might be ill-suited for applications in model monitoring and oversight.  [^1] Code and data available at: https://github.com/AI4LIFE-GROUP/sae_robustness.
Evaluating Adversarial Robustness of Concept
Representations in Sparse Autoencoders
Aaron J. Li1, Suraj Srinivas2, Usha Bhalla3, Himabindu Lakkaraju3 1 University of California, Berkeley 2 Bosch Research 3 Harvard University 1aaronjli@berkeley.edu 2suuraj.srinivas@gmail.com 3usha_bhalla@g.harvard.edu, hlakkaraju@seas.harvard.edu
1 Introduction
@@
In this work, we propose incorporating robustness as a core component of the SAE evaluation agenda, and concept-based explanations in general. We begin by defining SAE robustness under input-level perturbations, followed by the specification of adversarial settings for a comprehensive robustness evaluation. Each setting is characterized by two user-defined goals which determine how the input-level attack is applied. First, the semantic goal: the perturbation could be either untargeted, aiming to alter the SAE activations from the original, or targeted, aiming to match the perturbed activations to that of another semantically unrelated input. Second, the activation goal: either at the population level, where the objective is to manipulate all SAE neurons simultaneously; or at the individual level, where the goal is to activate a specific SAE feature. Putting these together, we define four different scenarios to evaluate SAE concept robustness against adversarial perturbations.
To search for adversarial inputs that manipulate SAE interpretations, we adapt Greedy Coordinate Gradient (GCG) (Zou et al., 2023), a widely used input-space search algorithm originally developed to generate adversarial prompts that induce harmful outputs in LLMs, to our SAE settings. An example of our attack is shown in Figure 1. Our findings reveal that current SAEs are not sufficiently robust to serve as reliable tools for interpreting and monitoring LLM behaviors, nor as trustworthy components in downstream applications that depend on stable, concept-level representations. We further believe that this vulnerability to adversarial input manipulation could generalize to other SAE-inspired approaches such as transcoders (Dunefsky et al., 2024) and crosscoders (Lindsey et al., 2024), highlighting the need for future research on developing more robust and reliable concept-extraction methods in the field of mechanistic interpretability.
Figure 1: An example of successful targeted population-level attack that doubles the concept overlap between $x_{1}$ and $x_{2}$ with only one adversarial token replacement.
@@
  • We propose a comprehensive and theoretically grounded evaluation framework that defines SAE robustness along semantic and activations goals.
  • We conduct extensive experiments by designing adversarial input-level attacks, showing that SAE interpretations are consistently vulnerable across multiple open-source LLMs, pretrained SAEs, and datasets.
2 Related Work
@@
3 Evaluating the Robustness of SAE Interpretations
@@ 3.1 Proposed Theory of SAE Robustness
$\exists c_{1},c_{1}^{\prime}\in\mathcal{C}:d_{c}(c_{1},c_{1}^{\prime})<\epsilon_{c},d_{z}(z_{1},z_{1}^{\prime})>\delta_{z}$
$\exists c_{1},c_{2}\in\mathcal{C}:d_{c}(c_{1},c_{2})>\delta_{c},d_{z}(z_{1},z_{2})<\epsilon_{z}$
@@
  • They support more fine-grained control, allowing perturbations in arbitrary directions and with variable step sizes in the input space (as explained in Section  3.2).
@@
$\max_{x_{1}^{\prime}}d_{z}(z_{1},z_{1}^{\prime})\quad\text{subject to}\hskip 5.0ptd_{x}(x_{1},x_{1}^{\prime})\leq\epsilon_{x}$ (3)
$\min_{x_{2}}d_{z}(z_{1},z_{2})\quad\text{subject to}\hskip 5.0ptd_{x}(x_{1},x_{2})\geq\delta_{x}$ (4)
@@ 3.2.1 Semantic Goal
$\min_{x_{1}^{\prime}}d_{z}(x_{1}^{\prime},x_{2})\quad\text{subject to}\hskip 5.0ptd_{x}(x_{1},x_{1}^{\prime})\leq\epsilon_{x}$ (5)
@@ 3.2.2 Activation Goal
The goal is to manipulate groups of SAE features simultaneously to control the overall sparse representation vector, so the distance metric can be defined by the overlap ratio between two sets of $k$ most activated SAE latents (we call this metric concept overlap ratio throughout this work):
@@
The perturbation can be directed toward a specific SAE feature by modifying its rank among all latent dimensions. In the case of activation, the objective is to increase the feature’s rank until it has non-zero activation after $\phi$. Conversely, in the case of deactivation, the objective is to reduce the feature’s rank such that the latent becomes zero. We thus define $d_{z}$ in a binary manner:
@@ 3.2.3 Perturbation Mode
To keep the input-level perturbations minimal, in this work our attacks are only allowed to change one token of the original input $x_{1}$. Specifically, for an input sequence with $n$ tokens, we launch $n$ independent attacks that each tries to adversarially replace one original token. Additionally, we also consider the attack that appends an one-token adversarial suffix to the input, so the attack performance to be reported would be the average across these $n+1$ individual trials.
@@ 3.3 Generalized Input-level Attack for SAE
The complete pseudocode for our generalized input-level attack is provided in Appendix B.
4 Experiments
In this section, we first describe our experimental setup, including model and dataset choices, followed by results across all evaluation settings. We then present several additional analyses to decouple SAE and LLM representational robustness, study non-adversarial SAE robustness, examine robustness trends across model depths, assess cross-model attack transferability, and finally conclude with a case study on manipulating well-annotated highly interpretable SAE latents.
4.1 Experiment Setup
Models
We evaluate robustness across open-source LLM-SAE pairs, including Gemma2-2B and Gemma2-9B from Gemmascope with JumpReLU SAEs varying in width, sparsity, and layer depth, as well as Llama3-8B with Top-K SAEs (Gao et al., 2024) as a separate model family. In this section, we present the results of layer 30 of Gemma2-9B (with an SAE width of 131K) in Figure 2, and more comprehensive evaluation results are included in Appendix D and Appendix G.
Datasets
We evaluate SAE robustness on four different datasets: AdvBench, AG News, SST2, and another generated dataset Art & Science. For each combination of dataset, LLM, and SAE, we select $200$ $(x_{1},x_{2})$ pairs with initial population-level SAE concept overlap ratio less than $0.30$ for robustness evaluation (for Gemma2-2B models we relax this threshold to $0.35$), so the evaluation data for different LLM-SAE pairs are not identical. We provide brief descriptions of the four datasets in Table 1. For the AdvBench dataset presented in Figure 2, the $x_{1}$ instances are harmful user instructions, while the $x_{2}$ instances are the same GPT-generated benign prompts used by Kumar et al. (2023).
Name $x_{1}$ $x_{2}$ # Tokens
AdvBench Harmful Benign 10 to 20
AG News Business Sports 5 to 10
SST2 Positive Negative 10 to 20
Art and Science Art/Humanities Science/Tech 20 to 30
Table 1: Descriptions of four datasets used in our evaluation, including the specific categories of $x_{1}$ and $x_{2}$ instances, as well as typical sequence lengths.
Evaluation Configurations
Since the residual streams of LLMs encode both semantic features and next-token prediction information, we append a short instruction prompt to the original sequence, "The previous sentence is about", to the input to better extract LLM’s semantic content from the last hidden state. At the individual level, we focus on $5$ SAE latents selected for activation or deactivation based on the semantic goal. For untargeted tasks, we choose the latents with the lowest or highest activation values for $x_{1}$, depending on the activation/deactivation setting. For targeted tasks, we select latents that are currently highly activated by $x_{2}$ but deactivated by $x_{1}$ (for the activation setting), or vice versa. The specific choices of hyperparameters are included in Table 5.
Figure 2: Attack Results of Gemma2-9B (131k) on AdvBench. Standard deviations are computed across 55 experiment runs.
4.2 Results Interpretation
We present the complete results of Gemma2-9B (131k) on the AdvBench dataset in Figure 2. For tasks on population level, we report concept overlap ratios before and after the attacks, while for the individual level we report the attack success rates (ASR). As mentioned in Section 3.2, each percentage value is averaged across $n+1$ independent attacks targeted on different token indices. More comprehensive evaluation results for other LLMs, SAEs, and datasets are included in Appendix D, which all demonstrate similar trends. In general, our attacks are effective in most cases, and we discuss several important insights below.
Successful Cases
Our population-level attacks are effective: a single-token adversarial perturbation can nearly double the concept overlap with an unrelated target sequence $x_{2}$ while reducing overlap with the original $x_{1}$ to roughly $20\%$$30\%$. We also provide an example of detailed comparison of top 5 activated SAE latents between a pair of $x_{1}$ and $x^{\prime}_{1}$ in Figure 3. At the individual level, over $60\%$ of the selected SAE latents are manipulable, and in the deactivation setting this fraction rises to about $90\%$.
Failure Cases
The adversarial attack only consistently fails on the untargeted tasks at the individual level: the SAE latents with lowest activation values cannot be activated by input perturbations. This is most likely due to dead latents (i.e. latents cannot be activated under any circumstances) prevalent in pretrained SAEs, which is aligned with previous findings by Gao et al. (2024).
Untargeted vs. Targeted
Figure 3: An example of changes in top 5 activated SAE concepts between the original $x_{1}$ and the perturbed $x^{\prime}_{1}$: only 2 out of the 5 original concepts remain on top. Top concepts in $x_{1}$ have cold colors, while new concepts introduced by $x^{\prime}_{1}$ have warm colors. The natural language annotations of the SAE latents are provided by Neuronpedia.
4.3 Additional Analyses
Decoupling SAE and LLM Robustness
To disentangle the robustness of the SAE from that of the underlying LLM, we analyze whether high-level semantic concepts can still be extracted from the perturbed $x^{\prime}_{1}$. First, we train a linear probe on top of LLM activations to detect a dataset-level concept via binary classification, using held-out data for each (dataset, model) pair. The four detected concepts are: (1) whether the prompt is harmful, (2) whether a news title is related to business, (3) whether the prompt has positive sentiment, and (4) whether the prompt relates to art or the humanities. All linear probes achieve at least $90\%$ test accuracy on the corresponding set of original $x_{1}$ instances. Then, we sample $600$ perturbed examples $x^{\prime}_{1}$ per probe and use it to detect the target semantic concept. Because the adversarial attacks produced many more candidates, we randomly subsample $600$ to ensure balanced comparisons across SAE widths and other evaluation settings. We call the fraction of prompts that the probe doesn’t detect the semantic concept on them as the False Negative Rate (FNR), and report the differences $\Delta\text{FNR}=\text{FNR}(x^{\prime}_{1})-\text{FNR}(x_{1})$ in Table 2. The lack of meaningful differences indicates that the dataset-level semantic concepts remain largely preserved in the LLM activations after adversarial perturbation, suggesting our attacks do not substantially alter the base model’s internal representations. As an alternative qualitative check, we manually review model self-summarizations of $x_{1}$ and $x^{\prime}_{1}$ using the short prompt mentioned in Section 4.1. We provide some examples in Figure 14. Although generations sometimes differ (perhaps due to randomness), adversarial tokens are rarely reflected in these self-summarizations.
Model Dataset $\Delta\text{FNR(\%)}$
Gemma2-2B AdvBench +2.3+2.3
AG News +1.8+1.8
SST2 +4.5+4.5
Art & Science +1.0+1.0
Gemma2-9B AdvBench +0.8+0.8
AG News +0.4+0.4
SST2 +1.5+1.5
Art & Science $0.0$
Llama3-8B AdvBench $0.0$
AG News $0.0$
SST2 +2.3+2.3
Art & Science $0.0$
Table 2: The differences in false negative rates between the original $x_{1}$ and the perturbed $x^{\prime}_{1}$ are not significant.
Constrained Synonym Attack
Instead of optimizing the GCG and evaluation losses in search of adversarial tokens, we are also interested in the SAEs’ generic robustness property when the input-level perturbations are not necessarily adversarial. To investigate this, we simply query an external LLM (GPT-4.1-mini) to generate $50$ different perturbed inputs with one token replaced by an synonym. The specific indices will be determined by the LLM, which is instructed to be diverse. Then we find the average attack performance achieved by the $50$ $x^{\prime}_{1}$ instances based on the same evaluation metrics used for adversarial attacks. Note that we focus only on untargeted synonym attacks, since they better reflect common real-world use cases. Based on Table 3, replacing a word with a synonym reduces the concept-overlap ratio by roughly $30\%$, indicating that SAE concept representations are not robust at the population level even under non-adversarial perturbations. These results also corroborate the effectiveness of our adversarial attacks: in the targeted setting, the best overlap achieved between $x^{\prime}_{1}$ and $x_{2}$ is only about $15\%$ (in absolute value) lower than the overlap obtainable via a single-word synonym replacement.
Model & SAE Population Individual
Activation Deactivation
Gemma2-2B (16k) $73.4$ $0.0$ $0.0$
Gemma2-2B (65k) $70.9$ $0.0$ $0.0$
Gemma2-9B (16k) $71.5$ $0.0$ $0.0$
Gemma2-9B (131k) $68.7$ $0.0$ $0.0$
Llama3-8B (131k) $66.2$ $0.0$ $0.0$
Table 3: Results of constrained synonym: a single-word synonym replacement can lead to about $30\%$ decrease in concept overlap, while it cannot change the individual activation status of selected SAE latents.
Robustness Change Across Model Depth
The main results presented in this section focus on a single mid-to-late layer for each model. To assess the generalizability of our findings across different model depths, we apply our population-level attacks to additional layers in Gemma2-9B and Llama3-8B (both with SAE widths of 131k). Empirically, we find that our attacks remain effective across layers, despite slight decreases in performance upper bound. The experiment results and relevant discussions are included in Appendix G.
Transferability of Attacks
In reality, the adversary might want to craft a single perturbed sequence that could be used to attack different models, so it’s important to investigate the transferability of our attacks across different LLMs. Since SAE latents of different models encode completely different semantic concepts, we only investigate transferability at the population level. As presented in Figure 16, although we observe notable performance degradations after model transfer, the attacks are still effective in matching target $x_{2}$ and perturbing original $x_{1}$.
Deactivating Highly Interpretable SAE Latents from Neuronpedia
In the above experiments, the set of manipulated SAE latents is empirically determined by our datasets (i.e., the most or least activated ones by our inputs). In fact, such evaluation procedure could be reversed: we may instead begin by selecting specific target SAE latents and then assess their robustness by identifying input sequences that strongly activate them. To explore this, we select several SAE latents associated with consistent and highly interpretable semantic concepts from Neuronpedia, and apply the untargeted individual-level attacks to top-activating sequences drawn from an external text corpus. Through experiments, we find that our attack could successfully deactivate these highly meaningful SAE latents. Illustrative examples are provided in Appendix I.
5 Discussions
Across extensive experiments and analyses, we find that SAE concept representations are not robust in most of the settings we have evaluated. However, this should not be read as a verdict against SAEs: an important future direction would be developing novel methods to denoise SAE activations and isolate the reliable signals. In this context, the attacks we introduce in this work might help identify and filter out non-robust SAE interpretations.
Meanwhile, we consider our evaluation framework and adversarial attacks as a general methodology for assessing concept-extraction tools for LLMs. While we focus on standard SAEs in this work, the same vulnerabilities likely extend to other variants such as transcoders (Dunefsky et al., 2024) and crosscoders (Lindsey et al., 2024), which similarly lack structural constraints or robustness-aware objectives during training. We leave the exploration of such extensions to future work.
6 Conclusion
In this work, we investigate the robustness of SAEs under input-level adversarial perturbations and introduce a comprehensive evaluation framework spanning semantic and activation-level goals. Our experiments show that SAE interpretations are highly vulnerable to minimal input changes, even when the underlying LLM remains semantically stable, raising concerns about their reliability in realistic settings. To advance trustworthy interpretability, we hope our work motivates the development of more robust tools for understanding LLMs, as stability under real-world conditions is essential for aligning model behavior with human expectations.
Limitations
As discussed in section 4.1, we target medium-sized LLMs because smaller models, along with their SAEs, lack the capacity to distinguish semantically unrelated inputs. However, this choice results in slower attacks. In practical applications, users may only need to evaluate the most relevant scenario among these proposed settings. Additionally, future work could explore optimizing the attack pipeline to better balance effectiveness and efficiency.
Acknowledgement
This work is supported in part by the NSF awards IIS-2008461, IIS-2040989, IIS-2238714, AI2050 Early Career Fellowship by Schmidt Sciences, and research awards from Google, JP Morgan, Amazon, Adobe, Harvard Data Science Initiative, and the Digital, Data, and Design (D3) Institute at Harvard. The views expressed here are those of the authors and do not reflect the official policy or position of the funding agencies.
Appendix A Loss Functions Used in Different Evaluation Settings
We include the specific GCG loss functions and evaluation metrics in Table 4.
@@ Appendix A Loss Functions Used in Different Evaluation Settings
Appendix B Pseudocode for Adversarial Input-Level Attack
The detailed pseudocode for our generalized GCG attack for SAEs is provided as Algorithm 1.